Solve the equation
9y+5-2y= 2y+5-y+3

Answer by stanbon(26275)

Solve the equation
9y+5-2y= 2y+5-y+3
---
7y + 5 = y + 8
6y = 3
y = 1/2
============
Cheers,
Stan H.